Unit 5 Lab--Principles of Accounting II

Exercise 25-17

The South Division of Swifty Company reported the following data for the current year.

Sales $2,950,000

Variable costs 1,947,000

Controllable fixed costs 605,000

Average operating assets 5,000,000

Top management is unhappy with the investment center’s return on investment (ROI). It asks the manager of the South Division to submit plans to improve ROI in the next year. The manager believes it is feasible to consider the following independent courses of action.

1. Increase sales by $300,000 with no change in the contribution margin percentage.

2. Reduce variable costs by $155,000.

3. Reduce average operating assets by 5%.

(a) Compute the return on investment (ROI) for the current year

Return on Investment

%

(b) Using the ROI formula, compute the ROI under each of the proposed courses of action

Return on investment

Action 1

%
Action 2

%
Action 3

%
Answer #1

Controllable margin = 2950000-1947000-605000 = 398000

a) ROI = Controllable margin/Average operating assets = 398000/5000000 = 7.96%

b) Calculate ROI under following situation

Return on investment

Action 1 (300000*34%+398000)/5000000

10

%
Action 2 (398000+155000)/5000000

11.06

%
Action 3 398000/4750000

8.38

%